OpenStudy (anonymous):

can you help me solve 6b+4=2(2b+1) Please ! i really dont understand thiss!

OpenStudy (anonymous):

first off calm down, and it just sort of a balancing act. First off let me work it and show you

OpenStudy (anonymous):

do you know what b equals?

OpenStudy (anonymous):

no so that is the goal to understand it

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ok sorry that probably confused you more, but 6B+4=2(2B+1) I did 6b+2b is 8B 4+1is 5 multiplied by 2 is 10 6B+4=2(2B+1) Whatever you add, subtract multiply or divide on one side you have to do to the other side. but first let's multiply out 2(2B+1) 2(2B+1)=4B+2 because we multiply 2*2B and 2*1 6B+4=2(2B+1) 6B+4=4B+2 now subtract 4B from both sides 2B+4=2 now subtract 4 from both sides 2B=2-4 2B=-2 divide both sides by 2 B=-1 now go to this website and they have an algebra solver http://www.algebra.com/algebra/homework/Expressions-with-variables/Expressions-with-variables.faq.question.245858.html